OpenStudy (anonymous):

Which improper fraction is equivalent to 2 3/9 14/9 20/9 21/9 23/9

OpenStudy (anonymous):

2 3/9 = 2*9 + 3 / 9 = 21/9 is correct choice

OpenStudy (anonymous):

tnx